Find the slope of a line perpendicular to the line that passes through the following points (3,9) and (7,15)
Answer by checkley77(12844)

(3,9) and (7,15)
m=(15-9)/(7-3)
m=6/4
m=3/2 ans. for this slope.
Any perpendicular line will have a slope=-2/3